# VICtoR: Learning Hierarchical Vision-Instruction Correlation Rewards for Long-horizon Manipulation

### v1

@@ A.1 Qualitative Analysis
A.2 Accuracy of Task Knowledge Generation
@@
Appendix B Detailed Settings for the Experiment
B.1 Details about our simulated environment and tasks
@@
Our environment was developed using Coppeliasim [52], with Pyrep [53] serving as the coding interface, as shown in Figure 3. In this environment, there are four interactable objects: a light, a drawer, a blue block, and a box, each associated with different stages as detailed in B.1. The tested tasks can be any permutation of these stages.
@@
To train the reward models for VICtoR and other baselines, we collected a dataset in our environment. Note that VICtoR requires only motion-level data for training, while other baselines need full-length videos. Therefore, the data used to train VICtoR and other baselines differ in type but are comparable in total amount per motion or stage. For VICtoR, we provide only the motion-level demos, with each motion accompanied by 100 demos and corresponding motion descriptions. For training other baseline models, such as LIV [16] and LOReL [50], we provide the full-length demo videos, each paired with task descriptions and consisting of 100 demos per task.
B.2 Details of Stage Detector and Training Details of Motion Progress Evaluator
We provide details of the Stage Detector and Training Details of the Motion Progress Evaluator in Table 6 and Table 7. For the hyperparameter selection of $\lambda_{1,2,3}$, we tried various combinations but found that the model is not sensitive to these hyperparameters. Therefore, we have chosen to set them all to the same values.
@@ B.2 Details of Stage Detector and Training Details of Motion Progress Evaluator
B.3 Details of MPE’s Inference and Reward Formulation
@@
B.4 Training Details of Baseline Models
@@
B.5 Training Details of Policy Training
Training details.
We trained the policy under the SubprocVecEnv from stable-baseline3 [55] with reward normalization by VecNormalize. For online interactable environment, we built a simulation environment with Coppeliasim [52] and Pyrep [53] and wrapped the environment for policy learning under the framework of OpenAI Gym [56]. Training time is about 3 hours for 10000 steps. Training Details of training of each task are listed in Table 5.
Table 10: Details of policy training
Policy PPO [51]
Learning Rate 0.003
Policy Model MlpPolicy
Discount factor ($\gamma$) 0.99
Reward Normalization True
Appendix C Learning Curves of Policy Learning
@@
Appendix D Additional Comparisons between VICtoR and Other Solutions to Robotic Policy Learning
@@
Appendix E More Potential Visualization for Composed Tasks in XSkill
@@
Appendix F Prompts for Task Knowledge Generation

### v2

@@ B.1 Qualitative Analysis - Misordered Action
B.2 Accuracy of Task Knowledge Generation
@@
B.3 Ablation Study on Three Contrastive Losses
To examine the importance and effectiveness of the three proposed contrastive objectives in Eq. 4, we train the Motion Progress Evaluator under three conditionsIn each experiment, we remove one contrastive loss, as illustrated in Table 4. This study is conducted on the “open box then pick blue block" task. The results clearly demonstrate that each objective plays a significant role, particularly the $L_{lfcn}$ objective. This objective effectively differentiates the progress in motion, providing nuanced guidance for long-horizon tasks.
Table 4: Ablation Study on the proposed three contrastive losses
Ablation Settings Success Rate
$\mathcal{L}_{tcn}$ $\mathcal{L}_{mcn}$ $\mathcal{L}_{lfcn}$
- 0.0%
- 61.8%
- 94.0%
96.4%
B.4 Ablation Study on Automatic Annotation by LVLMs
Since object status labels are required to deploy VICtoR in a new environment, we conduct an ablation study to automatically annotate these labels using modern LVLMs. The study is performed on the XSkill dataset, focusing on four interactable objects: cloth, light, oven, and drawer. Specifically, for each video where the robot interacts with one of these objects, we first use MDETR (without fine-tuning) to obtain cropped images of the target object. These cropped images are then provided to GPT-4V, which annotates the object’s status. Finally, we compare the consistency between human-annotated and LVLM-annotated labels, as summarized in Figure 9.
The results demonstrate a high level of consistency between the two sets of labels, indicating that deploying VICtoR in new environments is highly feasible. Most inconsistent labels arise during fuzzy phases when switching motions. These inconsistencies are easily corrected or even negligible.
Figure 9: Automatic Annotation by LVLMs: We leverage LVLMs (i.e., GPT-4V) to automatically annotate the status of objects in the XSkill dataset. The results are then compared with human-annotated labels to evaluate label consistency. The findings indicate that modern LVLMs already possess the capability to handle this task effectively. Most inconsistencies occur during the fuzzy phases when switching motions.
Appendix C Detailed Settings for the Experiment
C.1 Details about our simulated environment and tasks
@@
Our environment was developed using Coppeliasim [53], with Pyrep [54] serving as the coding interface, as shown in Figure 3. In this environment, there are four interactable objects: a light, a drawer, a blue block, and a box, each associated with different stages as detailed in Section C.1. The tested tasks can be any permutation of these stages.
@@
To train the reward models for VICtoR and other baselines, we collected a dataset in our environment. Note that VICtoR requires only motion-level data for training, while other baselines need full-length videos. Therefore, the data used to train VICtoR and other baselines differ in type but are comparable in total amount per motion or stage. For VICtoR, we provide only the motion-level demos, with each motion accompanied by 150 demos and corresponding motion descriptions. For training other baseline models, such as LIV [16] and LOReL [51], we provide the full-length demo videos, each paired with task descriptions and consisting of 150 demos per task.
C.2 Details of Stage Detector and Training Details of Motion Progress Evaluator
The structure of classifier $P$ for Stage Detector includes a CLIP Text Encoder for processing object state descriptions and a CLIP Image Encoder for processing cropped images. These encoders are followed by a downstream MLP classifier that scores relevance.
We provide more details of the Stage Detector and Training Details of the Motion Progress Evaluator in Table 7 and Table 8. For the hyperparameter selection of $\lambda_{1,2,3}$, we tried various combinations but found that the model is not sensitive to these hyperparameters. Therefore, we have chosen to set them all to the same values.
@@ C.2 Details of Stage Detector and Training Details of Motion Progress Evaluator
C.3 Details of MPE’s Inference and Reward Formulation
@@
C.4 Training Details of Policy Training
Training details.
We trained the policy under the SubprocVecEnv from stable-baseline3 [56] with reward normalization by VecNormalize. For online interactable environment, we built a simulation environment with Coppeliasim [53] and Pyrep [54] and wrapped the environment for policy learning under the framework of OpenAI Gym [57]. Training time is about 3 hours for 10000 steps. Training details for each task and policy are summarized in Tables 6 and 10, respectively.
Table 10: Details of policy training
Policy PPO [52]
Learning Rate 0.003
Policy Model MlpPolicy
Discount factor ($\gamma$) 0.99
Reward Normalization True
C.5 Computational cost and Resource requirements
The computational cost in our framework is 2 GPU hours for the training of the progress evaluator. For the policy learning, training with our reward model for 20000 steps required approximately 12 GPU hours. Both of the stats above are running on RTX 4090.
Regarding the inference time of the current implementation, the object detector takes approximately 51.5ms, and the reward model accounts for 24ms on average over 1000 steps (tested on a CPU Intel(R) Core(TM) i7-13700 and a GPU RTX 3090). Considering that end effector action latency is typically long in both simulation and real-world scenarios, we believe that a 0.05s inference time would not significantly impact the agent’s efficiency, even in a real-world environment.
C.6 Training Details of Baseline Models
@@
C.7 Details of experiments on XSkill
For the visualization experiments on real-world dataset XSkill [17] in Appendix F, we train VICtoR on 120 demonstrations with each breaking down into 3 different motion-level videos while baseline models are trained with the full-length demonstrations. Details about the 4 motion level skills and the test tasks in the visualization experiments are listed in Table 12.
Table 12: Text description for motions and test tasks in XSkill
Motion Name Description
close the drawer push the drawer so that it is closed
open the oven door open the oven door
open the light push the light on
move the cloth into the sink grab the cloth and release it into the sink
Test Tasks
close the drawer then move the cloth into the sink then open the light
close the drawer then move the cloth into the sink then open the oven door
close the drawer then open the light then move the cloth into the sink
close the drawer then open the light then open the oven door
open the light then move the cloth into the sink then open the oven door
open the light then close the drawer then move the cloth into the sink
open the light then open the oven door then close the drawer
open the oven door then open the light then move the cloth into the sink
Appendix D Learning Curves of Policy Learning
@@
Appendix E Additional Comparisons between VICtoR and Other Solutions to Robotic Policy Learning
@@
Motion Decomposition
Previous studies [41; 58] have utilized motion decomposition to facilitate learning in long-horizon manipulation tasks. By breaking down these tasks into subgoals, they demonstrated that imitation policy networks could be trained more effectively. While we also use the concept of task decomposition, our method introduces a tailored framework for VIC reward models, as shown in Figure 1. We segment tasks into three levels: stage, motion, and progress, providing a finer-grained approach to handle task complexities. Our experiments show that this structured method significantly improves the effectiveness of VIC reward models in complex scenarios.
@@
Appendix F More Potential Visualization for Composed Tasks in XSkill [17]
@@
Appendix G Prompts for Task Knowledge Generation